1. Plot the following Price and Quantity combinations: (4, 8), (1, 2), (5, 10) 2. Is your graph more likely to be a demand curve or a supply curve? Why? 3. Using the equation of a line, and P for price and Q for quantity, what is the algebraic formula of this curve?


1
Expert's answer
2022-03-02T10:25:40-0500

1. Plot the following Price and Quantity combinations: (4, 8), (1, 2), (5, 10)



2. Is your graph more likely to be a demand curve or a supply curve? Why?

It is a supply curve. Because the market supply curve is an upward sloping curve depicting the positive relationship between price and quantity supplied.


3. Using the equation of a line, and P for price and Q for quantity, what is the algebraic formula of this curve?

we will use two points to get the formula equation of the curve:

We will use (4, 8), (1, 2) coordinates.


"\mathrm{Find\:the\:line\:}\mathbf{y=mx+b}\mathrm{\:passing\:through\:}\left(4,\:8\right)\mathrm{,\:}\left(1,\:2\right)"


First, is to get the slope

"\mathrm{Slope}=\frac{y_2-y_1}{x_2-x_1}"


"\left(x_1,\:y_1\right)=\left(4,\:8\right),\:\left(x_2,\:y_2\right)=\left(1,\:2\right)"


"m=\frac{2-8}{1-4}"


Refine:

"m=2"

using y = mx + b

we can compute b as:

8 = 2 × 4 + b

b = 8 - 8 = 0

Therefore, b = 0,

y = 2x + 0 = 2x


Thus, the algebraic formula of this curve is "y = 2x"